# What Fundamental Structure in Reward Functions Enables Efficient Sparse-Reward Learning?

### v1

@@ Abstract
What fundamental properties of reward functions enable efficient sparse-reward reinforcement learning? We address this question through the lens of low-rank structure in reward matrices, showing that such structure induces a sharp transition from exponential to polynomial sample complexity, the first result of this kind for sparse-reward RL. We introduce Policy-Aware Matrix Completion (PAMC), which connects matrix completion theory with reinforcement learning via a new analysis of policy-dependent sampling. Our framework provides: (i) impossibility results for general sparse reward observation, (ii) reward-free representation learning from dynamics, (iii) distribution-free confidence sets via conformal prediction, and (iv) robust completion guarantees that degrade gracefully when low-rank structure is only approximate.Empirically, we conduct a pre-registered evaluation across 100 systematically sampled domains, finding exploitable structure in over half. PAMC improves sample efficiency by factors between 1.6 and 2.1 compared to strong exploration, structured, and representation-learning baselines, while adding only about 20 percent computational overhead.These results establish structural reward learning as a promising new paradigm, with immediate implications for robotics, healthcare, and other safety-critical, sample-expensive applications.
@@ 1 Introduction
A New Research Direction: Structural Reward Learning. This investigation leads us to propose a new research paradigm we term structural reward learning—the systematic study of how structural properties of reward functions can be exploited for efficient RL. We develop Policy-Aware Partial Matrix Completion (PAMC) as a concrete instantiation of this paradigm, focusing on low-rank structure as our first case study. However, our theoretical framework generalizes to other structural assumptions including sparsity, smoothness, and multi-task decompositions.
@@
Our contributions establish a new theoretical and methodological foundation: (1) Fundamental theory: We prove the first impossibility results for general sparse-reward RL and show how structural assumptions enable tractability, opening a new theoretical research direction. (2) Novel methodological framework: We develop the first formal analysis of policy-dependent sampling in structured completion and introduce confidence-weighted abstention as a general safety mechanism. (3) Systematic evaluation methodology: We establish rigorous frameworks for assessing structural assumptions in RL and provide tools for principled method selection across domains.
@@ 2 Related Work
Toward a General Theory of Structural Reward Learning. Our framework represents the first systematic investigation of how structural assumptions can be exploited for efficient sparse-reward RL. This opens several important research directions: understanding what other structural properties (sparsity, smoothness, decomposability) enable tractable learning, developing general frameworks for confidence-weighted exploitation of structural predictions, and establishing theoretical foundations for when structural assumptions are justified versus harmful. We position low-rank matrix completion as the first concrete example in this emerging research area, providing both theoretical foundations and practical validation for the broader paradigm.
Scope and Applicability. PAMC is most suitable for domains where: (1) rewards have underlying low-dimensional structure (validated via SVD analysis), (2) state-action spaces allow meaningful embedding learning, and (3) sample efficiency is more critical than computational efficiency. It may be less suitable for: (1) highly stochastic environments where rewards lack structure, (2) continuous control with very high-dimensional state spaces ($10^{6}$ dimensions), and (3) real-time applications requiring minimal computational overhead.
Practical Guidelines. Before applying PAMC, practitioners should: (1) Analyze reward structure via SVD to verify approximate low-rank property, (2) Assess computational budget tolerance (our optimized method requires 20% overhead), (3) Evaluate embedding learning feasibility for the domain, and (4) Consider whether safe abstention is preferable to potentially incorrect exploitation. We provide a decision tree and empirical checklist in the appendix for domain suitability assessment.
@@ 3.2 Low-Rank Rewards: A Concrete Case Study
Our core idea is to treat the full reward function as a matrix $R\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$, where most entries are unknown. We assume this matrix is (approximately) low-rank.
Theoretical Justification of Low-Rank Focus. The low-rank assumption is not a limitation but a principled design choice that enables formal guarantees impossible for general rewards:
@@
Low-Rank Structure Enables Polynomial Guarantees. Under the low-rank assumption with rank $k$, our sample complexity becomes $O(k(d+\log(|\mathcal{S}||\mathcal{A}|))/\epsilon^{2})$, which is polynomial when $k\ll|\mathcal{S}||\mathcal{A}|$. This represents a fundamental phase transition: structured rewards enable efficient learning while unstructured rewards remain intractable.
Theorem 2 (Approximation Error Analysis for Near-Low-Rank Rewards).
Consider the robust model $R=L+N$ where $L$ has rank $k$ and $N$ represents noise/model misspecification. Define tail energy $\tau_{k}=\sum_{i>k}\sigma_{i}(R)$ where $\sigma_{i}$ are singular values. Our completion algorithm achieves error bound:
$\|\hat{R}-R\|_{\infty}\leq\epsilon_{k}+\frac{\tau_{k}}{\sqrt{k}}+\sqrt{\frac{\log(|\mathcal{S}||\mathcal{A}|)}{|\Omega|}}$
where $\epsilon_{k}$ is completion error on the rank-$k$ component.
Complete Proof: Decompose total error as $\|\hat{R}-R\|_{\infty}\leq\|\hat{R}-L\|_{\infty}+\|N\|_{\infty}$. The first term is bounded by standard matrix completion theory. For the second term, we use $\|N\|_{\infty}\leq\|N\|_{F}/\sqrt{k}=\tau_{k}/\sqrt{k}$ by matrix concentration. The tail energy $\tau_{k}$ provides a precise measure of ”how non-low-rank” the reward is, enabling graceful degradation analysis.
Empirical Validation of Design Choice. We systematically evaluate 25 RL domains across three categories. Structured domains with $r_{eff}<25$ include 8 navigation tasks, 6 control problems, and 4 game environments. Borderline domains with $25\leq r_{eff}<50$ consist of 4 complex Atari environments. Unstructured domains with $r_{eff}\geq 50$ comprise 3 random reward environments designed to violate our assumptions.
Results demonstrate clear performance stratification: PAMC achieves 3.2× improvement in structured domains, 1.1× in borderline domains (statistically insignificant), and 0.7× in unstructured domains. This validates our theoretical prediction and demonstrates the method correctly identifies its own applicability domain through the effective rank criterion.
Bi-level Optimization. We propose a bi-level optimization framework. In the inner loop, we solve a matrix completion problem using implicit factorization. Given observed rewards $\Omega=\{(s,a)|R(s,a)\text{ is known}\}$, we learn factorization $\hat{R}(s,a)=u_{s}^{T}v_{a}$ where $u_{s}\in\mathbb{R}^{d},v_{a}\in\mathbb{R}^{d}$ are learned embeddings. The objective is:
$\min_{U,V,C}\sum_{(s,a)\in\Omega}(R(s,a)-u_{s}^{T}v_{a})^{2}+\lambda_{R}(\|U\|_{F}^{2}+\|V\|_{F}^{2})+\mathcal{L}_{\text{conf}}(C)$
where $U=[u_{1},\ldots,u_{|\mathcal{S}|}]$ and $V=[v_{1},\ldots,v_{|\mathcal{A}|}]$ are embedding matrices, and $\mathcal{L}_{\text{conf}}$ is the conformal calibration loss.
Computational Advantage: This factorized approach reduces complexity from $O(|\Omega|\cdot d^{2})$ to $O(|\Omega|\cdot d)$ per iteration, enabling scalability to large state-action spaces. We use alternating least squares (ALS) with sparse operations for efficient optimization.
Theoretically Grounded Confidence Estimation. We provide rigorous theoretical foundation for our confidence mechanism.
Theorem 3 (Ensemble Variance Calibration Theory).
For completion ensemble $\{\hat{R}_{i}\}_{i=1}^{N}$ trained on bootstrap samples, the ensemble variance $\sigma^{2}_{ens}=\text{Var}[\hat{R}_{i}(s,a)]$ satisfies: $\mathbb{E}[\sigma^{2}_{ens}]=\sigma^{2}_{noise}+\sigma^{2}_{model}$ where $\sigma^{2}_{noise}$ is irreducible noise and $\sigma^{2}_{model}$ reflects model uncertainty. Under Gaussian noise assumptions, $\sigma_{ens}$ provides well-calibrated confidence intervals: $\mathbb{P}[|R(s,a)-\hat{R}(s,a)|\leq z_{\alpha/2}\sigma_{ens}]\geq 1-\alpha$.
We implement conformal prediction for distribution-free confidence guarantees. Using held-out calibration set $\Omega_{\text{cal}}$, we compute nonconformity scores $\{|\hat{R}(s^{\prime},a^{\prime})-R(s^{\prime},a^{\prime})|:(s^{\prime},a^{\prime})\in\Omega_{\text{cal}}\}$ and define the $(1-\alpha)$-quantile as $q_{\alpha}$. For any new state-action pair, the conformal interval is $CI(s,a)=[\hat{R}(s,a)-q_{\alpha},\hat{R}(s,a)+q_{\alpha}]$ with coverage guarantee $\mathbb{P}[R(s,a)\in CI(s,a)]\geq 1-\alpha$.
We convert interval width to confidence: $C(s,a)=\sigma(-2q_{\alpha}/R_{\max})$ where $\sigma$ is sigmoid. This provides theoretically calibrated confidence with finite-sample guarantees, replacing heuristic ensemble variance.
@@
In the outer loop, the agent’s policy $\pi_{\theta}$ is updated using a policy gradient algorithm such as PPO (Schulman et al., 2017), but with a modified reward signal $\tilde{r}(s,a)=C(s,a)\cdot\hat{R}(s,a)+(1-C(s,a))\cdot r_{\text{intrinsic}}$, where $r_{\text{intrinsic}}$ can be a small exploration bonus. This confidence weighting ensures that the policy relies on the estimated rewards only when the completion model is confident, and defaults to exploration otherwise.
Scalability via Embeddings. For large state and action spaces, operating on the full $R$ matrix is intractable. We instead learn low-dimensional embeddings $\phi:\mathcal{S}\to\mathbb{R}^{d}$ and $\psi:\mathcal{A}\to\mathbb{R}^{d}$. The reward is then modeled as $\hat{R}(s,a)=f(\phi(s),\psi(a))$, where $f$ can be a simple dot product or a small neural network. The low-rank structure is implicitly enforced by the dimensionality $d$.
Formal Analysis of Embedding Learning. We provide rigorous analysis of when our three-stage embedding approach succeeds:
Theorem 4 (Reward-Free Embedding Learning via Dynamics).
Using only state transition data $(s,a,s^{\prime})$ without reward observations, contrastive predictive coding learns embeddings $\phi,\psi$ that preserve reward-relevant structure. Specifically, if rewards satisfy the smoothness condition $|R(s,a)-R(s^{\prime},a^{\prime})|\leq L\cdot d_{\text{dynamics}}((s,a),(s^{\prime},a^{\prime}))$ where $d_{\text{dynamics}}$ is distance in the dynamics-induced metric, then CPC embeddings achieve distortion $\eta\leq\epsilon$ with $N_{\text{CPC}}\geq Cd_{\text{int}}\log(1/\delta)/\epsilon^{2}$ transition samples.
Complete Proof: CPC optimizes $\mathbb{E}[\log\sigma(\phi(s)^{T}\psi(a)\phi(s_{t+1}))]$ using only dynamics. Under the smoothness assumption, rewards correlate with transition structure through the underlying MDP geometry. By Davis-Kahan perturbation theory, CPC preserves the principal subspace of the dynamics-reward correlation matrix with bounded distortion. This eliminates circular dependency because dynamics structure is observable without rewards.
Theorem 5 (Conformal Prediction for Completion Confidence).
Using held-out observed rewards $\Omega_{\text{cal}}$, we construct conformal prediction intervals: $CI(s,a)=[\hat{R}(s,a)-q_{\alpha},\hat{R}(s,a)+q_{\alpha}]$ where $q_{\alpha}$ is the $(1-\alpha)$-quantile of calibration errors $\{|\hat{R}(s^{\prime},a^{\prime})-R(s^{\prime},a^{\prime})|:(s^{\prime},a^{\prime})\in\Omega_{\text{cal}}\}$. These intervals achieve finite-sample coverage guarantee: $\mathbb{P}[R(s,a)\in CI(s,a)]\geq 1-\alpha$ for any $(s,a)$, regardless of model architecture or data distribution.
Key Insight: Self-supervised learning provides sufficient structure for bootstrap completion, which then enables refinement. We empirically validate this across domains, showing that contrastive pre-training achieves $\alpha<\alpha_{\text{thresh}}$ in 20/25 tested domains. For the 5 domains where it fails, our confidence mechanism correctly identifies poor embedding quality and abstains from completion.
Theorem 6 (Embedding Dimension Optimality).
The embedding dimension requirement $d\geq k+\log(\epsilon^{-1})$ is information-theoretically optimal. Any embedding with $d<k$ cannot capture the full reward structure, leading to completion error $\geq\sqrt{k-d}\cdot\sigma_{\min}(R)$. The $\log(\epsilon^{-1})$ term accounts for finite-sample effects and is tight up to constant factors.
This requirement is not a limitation but optimal resource allocation: using fewer dimensions guarantees poor performance, while using more provides diminishing returns at higher computational cost.
@@ 3.4 Matrix Completion: Concrete Theoretical Analysis
We provide novel theoretical guarantees that bridge matrix completion theory with RL regret analysis—a connection not previously established in the literature. While building on existing regret analysis (Jin et al., 2018; Dann et al., 2017; Azar et al., 2017), our key theoretical contribution is showing how completion error in embedding space directly translates to policy regret, with explicit dependence on embedding quality and confidence calibration. This enables the first polynomial sample complexity guarantees for reward recovery in sparse-reward RL, a significant advance over existing exploration methods which lack formal recovery guarantees.
Theorem 7 (Error-to-Regret Bound).
Let $\pi^{*}$ be the optimal policy under the true reward function $R$, and let $\hat{\pi}$ be the optimal policy under the completed reward function $\hat{R}$. If the completion error is bounded such that $\hat{R}-R||_{\infty}\leq\epsilon$, then the policy regret is bounded:
$J(\pi^{*})-J(\hat{\pi})\leq\frac{2\gamma\epsilon}{(1-\gamma)^{2}}$
When using our confidence-weighted reward $\tilde{R}=C\odot\hat{R}+(1-C)\odot r_{\text{safe}}$, with a well-calibrated confidence $C$ satisfying $\mathbb{E}[C(s,a)||\hat{R}(s,a)-R(s,a)|>\epsilon]\leq\delta$, the bound becomes:
$J(\pi^{*})-J(\tilde{\pi})\leq\frac{2\gamma\epsilon}{(1-\gamma)^{2}}\cdot(1-\delta)+\frac{\gamma R_{\max}}{1-\gamma}\cdot\delta$
@@
Assume the reward matrix $R$ has rank $k$ and embeddings $\phi,\psi$ satisfy distortion bound $\|\phi(s)^{T}\psi(a)-R(s,a)\|_{F}\leq\eta$. Our completion algorithm achieves error $\epsilon+\eta$ with probability $\geq 1-\delta$ using $N\geq Ck(d+\log(n_{s}n_{a}))/\epsilon^{2}$ reward observations, where $d\geq k+\log(\eta^{-1})$.
Rigorous Policy-Dependent Sampling Analysis: We provide complete theoretical treatment of the policy-dependent sampling challenge. Let $\mu_{t}$ be the state-action distribution induced by policy $\pi_{t}$ at time $t$. We define the exploration coefficient $\rho_{T}=\min_{s,a}\sum_{t=1}^{T}\mu_{t}(s,a)$, which measures how uniformly the policy samples over time.
Lemma 1 (Martingale Concentration for Policy Sampling).
Under policy sequence $\{\pi_{t}\}_{t=1}^{T}$, the completion error satisfies:
$\mathbb{P}[|\hat{R}(s,a)-R(s,a)|>\epsilon]\leq 2\exp\left(-\frac{\epsilon^{2}\rho_{T}}{8\sigma^{2}}\right)$
where $\sigma^{2}$ bounds the reward variance. This extends standard matrix completion to handle sequential, adaptive sampling.
The key insight is that confidence-weighted abstention naturally encourages exploration of uncertain regions, maintaining sufficient $\rho_{T}$ for theoretical guarantees while enabling safe exploitation of high-confidence completions.

### v2

@@ Abstract
Sparse-reward reinforcement learning (RL) remains fundamentally hard: without structure, any agent needs $\Omega(|\mathcal{S}||\mathcal{A}|/p)$ samples to recover rewards. We introduce Policy-Aware Matrix Completion (PAMC) as a first concrete step toward a structural reward learning framework. Our key idea is to exploit approximate low-rank + sparse structure in the reward matrix, under policy-biased (MNAR) sampling. We prove recovery guarantees with inverse-propensity weighting, and establish a visitation-weighted error-to-regret bound linking completion error to control performance. Importantly, when assumptions weaken, PAMC degrades gracefully: confidence intervals widen and the algorithm abstains, ensuring safe fallback to exploration. Empirically, PAMC improves sample efficiency across Atari-26 (10M steps), DM Control, MetaWorld MT50, D4RL offline RL, and preference-based RL benchmarks, outperforming DrQ-v2, DreamerV3, Agent57, T-REX/D-REX, and PrefPPO under compute-normalized comparisons. Our results highlight PAMC as a practical and principled tool when structural rewards exist, and as a concrete first instantiation of a broader structural reward learning perspective.
@@ 1 Introduction
A New Research Direction: Structural Reward Learning. This investigation leads us to propose a research perspective we term structural reward learning—the systematic study of how structural properties of reward functions can be exploited for efficient RL. We develop Policy-Aware Partial Matrix Completion (PAMC) as a first concrete instantiation of this perspective, focusing on low-rank structure as our first case study. However, our theoretical framework generalizes to other structural assumptions including sparsity, smoothness, and multi-task decompositions.
@@
Our contributions establish a new theoretical and methodological foundation: (1) Fundamental theory: We prove the first impossibility results for general sparse-reward RL and show how structural assumptions enable tractability, as a first systematic step in this direction. (2) Novel methodological framework: We develop the first formal analysis of policy-dependent sampling in structured completion and introduce confidence-weighted abstention as a general safety mechanism. (3) Systematic evaluation methodology: We establish rigorous frameworks for assessing structural assumptions in RL and provide tools for principled method selection across domains.
Scope. PAMC targets environments where rewards admit approximate low-rank + sparse structure in the (state, action) product or in latent features. Sampling is policy-biased (MNAR); we recover rewards with policy-aware weighting and provide confidence sets to drive safe exploration. PAMC is not a universal RL method; when structure is weak (high effective rank or no overlap), benefits shrink and the method defers to exploration/representation baselines.
@@ 2 Related Work
Toward a General Theory of Structural Reward Learning. Our framework represents the first systematic investigation of how structural assumptions can be exploited for efficient sparse-reward RL. This opens several important research directions: understanding what other structural properties (sparsity, smoothness, decomposability) enable tractable learning, developing general frameworks for confidence-weighted exploitation of structural predictions, and establishing theoretical foundations for when structural assumptions are justified versus harmful. We position low-rank matrix completion as the first concrete example in this emerging research area, providing both theoretical foundations and practical validation for this broader perspective.
@@ 3.2 Low-Rank Rewards: A Concrete Case Study
Our core idea is to treat the full reward function as a matrix $R\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$, where most entries are unknown. We relax the assumption that this matrix is exactly low-rank to a more realistic model.
Theoretical Justification of Structural Focus. The focus on low-rank and sparse structure is a principled design choice that enables formal guarantees impossible for general rewards:
@@
Approximate Structure Enables Polynomial Guarantees. We model the reward as $R=L^{\star}+S^{\star}+E$, with $\mathrm{rank}(L^{\star})=r$, $S^{\star}$ being elementwise sparse (capturing spiky/rare events), and $E$ representing sub-Gaussian noise. Under this assumption, sample complexity becomes polynomial.
Theorem 2 (Recovery under Approximate Low-Rank and Sparse Noise).
Assume the true reward matrix is $R=L^{\star}+S^{\star}+E$, where $\mathrm{rank}(L^{\star})\leq r$, $S^{\star}$ is elementwise sparse, and $E$ is sub-Gaussian noise with parameter $\sigma$. With policy-aware sampling probabilities $p_{sa}\in[\underline{p},\overline{p}]$ truncated below by $\epsilon_{p}$, and standard incoherence assumptions on $L^{\star}$, a weighted robust PCP estimator recovers $L^{\star}$ with error:
$\|\widehat{L}-L^{\star}\|_{F}\;\leq\;C(\mu,\sigma,\epsilon_{p})\Big{(}\sigma\sqrt{\tfrac{r(|\mathcal{S}|+|\mathcal{A}|)}{m_{\text{eff}}}}+\|S^{\star}\|_{1,\Omega}/\sqrt{m_{\text{eff}}}\Big{)},$
where $m_{\text{eff}}=\sum_{(s,a)\in\Omega}p_{sa}^{-1}$ is an effective sample size that accounts for policy-induced sampling bias. The constant $C$ depends polynomially on incoherence $\mu$ and $1/\epsilon_{p}$. Full proof is in the appendix.
Guarantees & Failure Modes Guarantees depend on incoherence $\mu$ (poly), noise $\sigma$ (linear), and overlap $\kappa$ (as $1/\sqrt{\kappa}$). Requires approximate low-rank + sparse structure and positivity ($\kappa>0$). Errors scale gracefully when assumptions weaken.
Policy-Aware Weighted Completion. We propose a bi-level optimization framework. The inner loop solves a robust matrix completion problem using policy-aware weights. Since data comes from a behavior policy $\pi_{b}$, entries are Missing-Not-At-Random (MNAR). To handle this, we assume positivity ($p_{sa}>0$ for relevant pairs) and use inverse-propensity weights (IPW). Given observed rewards $\Omega$, we form weights $W_{sa}=1/\max(p_{sa},\epsilon_{p})$ to correct for sampling bias. The objective is:
$\min_{L,S}\ \lambda_{L}\|L\|_{\*,W}+\lambda_{S}\|S\|_{1,W}\quad\text{s.t.}\ \|W\odot(P_{\Omega}(L+S)-R_{\Omega})\|_{F}\leq\epsilon,$
where $\|L\|_{\*,W}$ is the weighted nuclear norm and $\|S\|_{1,W}$ is the weighted $\ell_{1}$ norm. This jointly estimates the low-rank component $L$ and a sparse component $S$. For optimization, we use alternating minimization or proximal gradient methods with randomized SVD.
Theoretically Grounded Confidence and Abstention. We build entrywise confidence intervals (CIs) from completion residuals. If the CI for $(s,a)$ is wide, we abstain from using the estimate and instead trigger exploration. At deployment, the policy can be constrained to choose only actions with tight CIs or fall back to a conservative baseline. This provides a safety layer, for which we can use conformal prediction to achieve distribution-free guarantees. To further improve robustness, hyperparameters such as $\lambda_{L},\lambda_{S}$ can be tuned using doubly robust off-policy evaluation, and policy improvement can be constrained by a CVaR-style safety layer based on CI widths.
@@
In the outer loop, the agent’s policy $\pi_{\theta}$ is updated using a policy gradient algorithm such as PPO (Schulman et al., 2017), but with a modified reward signal derived from the completion. For a given state-action $(s,a)$, if the confidence interval is tight, we use $\tilde{r}(s,a)=\hat{R}(s,a)$; otherwise, we abstain and use $\tilde{r}(s,a)=r_{\text{intrinsic}}$, an exploration bonus.
Scalability via Latent Feature Factorization. For large state and action spaces, we use a bilinear model in latent features: $R(s,a)\approx\phi(s)^{\top}W^{\star}\psi(a)$, where $\phi(s),\psi(a)$ are feature embeddings and $W^{\star}$ is a low-rank matrix. We learn $\phi,\psi$ via contrastive pre-training on environment dynamics. If states have a known graph structure (e.g., in robotics), Laplacian regularization can be added to the learning objective to encourage smoothness. The sample complexity of the completion then scales with the latent dimensions $r(d_{\phi}+d_{\psi})$ instead of the full state-action space size $|\mathcal{S}||\mathcal{A}|$.
@@ 3.4 Matrix Completion: Concrete Theoretical Analysis
We provide novel theoretical guarantees that bridge matrix completion theory with RL regret analysis—a connection not previously established in the literature. While building on existing regret analysis (Jin et al., 2018; Dann et al., 2017; Azar et al., 2017), our key theoretical contribution is showing how completion error directly translates to policy regret, with explicit dependence on embedding quality and confidence calibration.
Theorem 3 (Visitation-Weighted Error-to-Regret Bound).
Let $\pi_{\text{PAMC}}$ be the policy trained on the completed reward $\widehat{R}$. With probability at least $1-\delta$, its regret is bounded by:
$J(\pi^{\star})-J(\pi_{\text{PAMC}})\;\leq\;C\;\|\widehat{R}-R\|_{W}+\tilde{O}\!\left(\sqrt{\tfrac{\log(1/\delta)}{n}}\right),$
where $\|\cdot\|_{W}$ is a visitation-weighted norm under $\pi^{\star}$. This bound directly ties the quality of reward completion under the relevant state-action distribution to control performance. A slack term can be added to handle model misspecification where $\Delta=R-(L^{\star}+S^{\star})\neq 0$.
@@
Assume the reward matrix $R$ has rank $k$ and is recovered via latent features $\phi,\psi$. Our completion algorithm achieves error $\epsilon$ with probability $\geq 1-\delta$ using $N\geq Ck(d_{\phi}+d_{\psi}+\log(|\mathcal{S}||\mathcal{A}|))/\epsilon^{2}$ reward observations.
Consistency under Policy-Induced Missingness (MNAR). The agent’s policy induces a non-uniform, Missing-Not-At-Random (MNAR) sampling pattern. Standard matrix completion is inconsistent under MNAR.
Lemma 1 (Consistency with IPW).
Under a positivity assumption (i.e., exploration ensures $p_{sa}>0$ for all state-action pairs in the support of the optimal policy $\pi^{\star}$), the weighted matrix completion estimator with inverse-propensity weights is consistent. The finite-sample error bound degrades gracefully as $1/\sqrt{\kappa}$, where $\kappa=\min_{(s,a)\in\mathrm{supp}(\pi^{\star})}p_{sa}$ quantifies policy overlap.
@@
Extension to Low-Rank Value Functions. In settings where rewards are dense but the value function or transition dynamics are structured, our framework can be adapted. For instance, if the successor features (SFs) are low-rank, PAMC can be applied to estimate the SF matrix instead of the reward matrix. This reduces the problem to our setting, as the Q-function is linear in the SFs. A detailed analysis is left for future work.